0 1
Add a comment Improve this question Transcribed image text
Answer #1

For a first order reaction k = 1/t ln [a0/a]

a0 = 0.36M

a = 0.30M

t = 15 min

hence k = 1/15 ln [0.36/0.30] = 0.012 min-1
Answer #2

Alternative meathod: